OpenStudy (anonymous):

Solve: five less than twice what negative number squared is 157?

OpenStudy (anonymous):

2(-x) - 5 = 157

OpenStudy (anonymous):

oops, I missed the squared

OpenStudy (anonymous):

157=2x^2-5

OpenStudy (anonymous):

yeah, that :)

OpenStudy (anonymous):

162=2x^2 81=x^2 x=9